OpenStudy (anonymous):

Find all solutions of sin(z)=0?

OpenStudy (amistre64):

all of them? and is z the complex variable or just some random variable that is real?

OpenStudy (anonymous):

z= complex

OpenStudy (amistre64):

every odd iteration of pi makes sin(x) = 0 ..... but Z eh

OpenStudy (amistre64):

thought I erased that odd part lol

OpenStudy (angela210793):

z=2*k*pi

OpenStudy (anonymous):

z=2kpi? Where K is a real no.? :D

OpenStudy (angela210793):

\[k \in Z\]

OpenStudy (amistre64):

I dont know of any sin(z) =0 that exists in the complex part

OpenStudy (angela210793):

Idk how u may call them...Sorry...

OpenStudy (amistre64):

if we start out as the [0,inf) real line as our initial ray, then any tine a + 0i occurs we get sin=0

OpenStudy (amistre64):

so at 1+0i and -1+0i times whatever scalar you wanna used for the unit circle

OpenStudy (anonymous):

But a+0i isn't a complex number, So you're saying every integer is a complex number, aren't you?

OpenStudy (amistre64):

every integer is a complex number in the same sense that every integer is a rational number; because they are inherently a part of the set..

OpenStudy (amistre64):

the complex plane includes the real number line; so yes... all real numbers are a subset of the complex plane

OpenStudy (anonymous):

OK, if so, then the answer is just one, which is 0.

OpenStudy (angela210793):

\[R \subset complex no.s\]

OpenStudy (amistre64):

its only '1' if we make a unit circle in the complex plane and choose the +Reals as our initial ray; then rotate around and terminate everytime we hit the real line again...

OpenStudy (amistre64):

when Z = any complex number of the form: a+0i; we get sin(z) = 0

OpenStudy (anonymous):

But only integer which is of the form npi is 0

OpenStudy (anonymous):

So the answer should be just one

OpenStudy (anonymous):

But now what confuses me is, x-1=0 has a complex solution?

OpenStudy (amistre64):

Z is not retricted to "off of the real number line" so just sating '1' is useless

OpenStudy (angela210793):

Not 1....4 example...k=-1---> Z=2(-1)*pi=-2pi=angle 0 k=0--->Z=angle 0 k=1--->=2*1*pi=angle 0

OpenStudy (amistre64):

the coordinate system is like the reals; but you define Z as: z = (x + yi)

OpenStudy (amistre64):

as opposed to: point = (x,y)

OpenStudy (anonymous):

Oh, so a+bi is a complex number, so is a + 0i so is R+0i Thus the solution to this would be npi+0i, wouldn't it?

OpenStudy (amistre64):

exactly :)

OpenStudy (amistre64):

but to say npi means to say only those z values that are a real multiple of pi is wrong

OpenStudy (amistre64):

every point along the real number line is valid

OpenStudy (amistre64):

pi is a number of the real number line; its a measure of the diatmer of a circle to its circumference

OpenStudy (anonymous):

Is the polar form correct, bcuz you are saying sin(1) = 0?

OpenStudy (amistre64):

sin(1 +0i) = 0 at least :)

OpenStudy (anonymous):

sin(1+0i) = 0?

OpenStudy (amistre64):

sin(z) = 0 means at least: sin(a + bi) = 0

OpenStudy (angela210793):

sin(1)=0....? how come?

OpenStudy (amistre64):

sin(a + 0i) = 0 for any given a

OpenStudy (anonymous):

But polar form just gives you a way to express

OpenStudy (anonymous):

Not a complex plane

OpenStudy (amistre64):

polar form can be used to interpret the complex plane into a real plane;

OpenStudy (anonymous):

Hmm., check this out http://www.wolframalpha.com/input/?i=sin%281%2B0i%29

OpenStudy (amistre64):

it wold resemble the polar form yes...

OpenStudy (angela210793):

Don't take me wrong PLEASE....but....why do u explain a very easy problem in this way?....O.o....ur making it harder then it really is....O.o

OpenStudy (angela210793):

than*

OpenStudy (anonymous):

Don't you think, we are leaving out someone in this discussion? The asker!

OpenStudy (amistre64):

i havent blocked them lol

OpenStudy (amistre64):

wolfram is good but still has issues with interpreting data

OpenStudy (angela210793):

haha lol :D.....maybe he has got it...:)

OpenStudy (anonymous):

hey..i'm still here,get confuse @_@

OpenStudy (angela210793):

Oops....@Amistre do u agree with me that the answer is Z=2kpi?

OpenStudy (amistre64):

im open to me being wrong :) but I would have to see how sin(z) is sposed to be interpreted :)

OpenStudy (angela210793):

K belongs to Z

OpenStudy (anonymous):

i'm reading this post, try to digest, what do u guys think: http://answers.yahoo.com/question/index?qid=20080529224140AAGdnOg

OpenStudy (amistre64):

angela, I would have to research it to tell :)

OpenStudy (angela210793):

Ok.....I have always thought math is same everywhere...but...o.O

OpenStudy (amistre64):

maybe sin(z) = 0 where e^(ix) = e^(-ix) :)

OpenStudy (angela210793):

maybe....ur the smart guy here anyway :P

OpenStudy (angela210793):

u have Z={..........-3,-2,-1,0,1,2,3.........} right?

OpenStudy (amistre64):

no; z is beig used in this context to indicate the complex variable and not the set of integers

OpenStudy (amistre64):

its like using x for a real variable; they use z for the complex variable

OpenStudy (angela210793):

Really??????? well in my answer i meant for this kind of Z.....and wht do u call ''my Z''?

OpenStudy (amistre64):

but all signs point to the answer being npi where n is any integer

OpenStudy (anonymous):

Integers!

OpenStudy (amistre64):

your Z would indicate "any integer" ;

OpenStudy (angela210793):

ahaa...that's wht i meant....:S:S

OpenStudy (angela210793):

Thanks :):):) ^_^

OpenStudy (anonymous):

Using 0 = (e^(iz) - e^(-iz)) / 2i I get stucked when 2iz= log 1 what's next...

OpenStudy (anonymous):

@angela210793: You never seemed to be not knowing that though! lol

OpenStudy (amistre64):

http://jwbales.us/precal/part6/part6.3.html this is more my thinking

OpenStudy (amistre64):

sin(z) = bi/|z|

OpenStudy (amistre64):

when bi=0; sin(z) = 0

OpenStudy (angela210793):

Hey...X( X( I know what that is X( is just that i thought integers in english was Z as in my place is...

OpenStudy (amistre64):

I saw one textbook have it as J and I was like... whats that? lol

OpenStudy (anonymous):

Why can't they keep it to just one universal symbol~!

OpenStudy (angela210793):

37 mins ago I've written:''Idk how u may call them...Sorry...'' So it's not my fault..:(:(:(

OpenStudy (anonymous):

ho ho ho

OpenStudy (amistre64):

i blame the chinese... cant trust anyone theat squints all the time..shifty beady little eyes lol

OpenStudy (angela210793):

hahahaha :D :D Lool :D :)

OpenStudy (anonymous):

\[z=k\pi, k\in Z\]

OpenStudy (anonymous):

what is with the latex???

OpenStudy (angela210793):

O.o

OpenStudy (amistre64):

its new since they upgraded today; nothing works lol,

OpenStudy (anonymous):

@satellite73, are u confident with that answer? which method u're using?